Vector a has magnitude of 8 units and makes an angle of 45 with the positive x-axis vector b also has the same magnitude of 8 units and directed along the negative x-axis find the magnitude and direction of A+B


1
Expert's answer
2021-07-05T08:42:17-0400
"\bold{A+B}=(8\cos{45},8\sin{45})+(-8,0)\\=(-2.3,5.7)\\A+B=6.1\\\theta=112.5\degree"